# Distributed Algorithms for Potential Problems

### v3

@@ 4 Analysis of the algorithm
This implies that there is a $j^{\star}\leq i-1$ such that $A$ is not contained in $\mathcal{N}_{t_{2}}[B_{j^{\star}}]$. Note that the diameter of $A$ is at most $c_{3}\operatorname{log}n/\varepsilon$ and at least one node of $A$ is not in $\mathcal{N}_{t_{2}}[B_{j^{\star}}]$. As a result, the set $\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}+2r+1}[A]$ is fully contained in some cluster $C^{(j^{\star})}_{s}$, for some $s\in[h_{j^{\star}}]$.
@@
$\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]\cap\left(\cup_{t\in[t_{i}]}\hat{S}_{t}\right),$
and such that $\operatorname{IR}(A^{\prime},\ell_{j^{\star}},\ell_{A^{\prime}})\geq R_{j^{\star}+1}-2\varepsilon$. By Lemma 4.5, for each $v\in A^{\prime}$, there exists a minimal improving set $(A^{\prime\prime},\ell_{A^{\prime\prime}})$ such that $A^{\prime\prime}\subseteq A^{\prime}\cap\mathcal{N}_{c_{3}\operatorname{log}n/\varepsilon}[v]$ and $\operatorname{IR}(A^{\prime\prime},\ell_{j^{\star}},\ell_{A^{\prime\prime}})\geq R_{j^{\star}+1}-3\varepsilon\geq R_{j^{\star}}$. The fact that
$A^{\prime\prime}\subseteq\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]$
implies that $A^{\prime\prime}\subseteq\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[A]$. Since $\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}+2r+1}[A]$ is fully contained in some cluster $C^{(j^{\star})}_{s}$, so it is $\mathcal{N}_{2r+1}[A^{\prime\prime}]$. Hence, the sequence $(S^{(j^{\star})}_{s,t},\ell_{S^{(j^{\star})}_{s,t}})_{t\in[k_{s}^{(j^{\star})}]}$ was not maximal, reaching a contradiction. ∎
@@
By Lemma 4.2, we have that the algorithm terminates. As discussed in Section 3.2, the running time of the algorithm is $O(\operatorname{log}^{6}n)$. We just have to argue about correctness.
Let $i=c_{1}\operatorname{log}n$ be the last phase, and let $\ell_{i}$ be the labeling obtained at the end of phase $i$ (Lines 18 and 23). By contradiction, suppose that there is an error in the graph, that is, a centered graph $(H,v_{H})$ of radius $2r$ that does not belong to $\mathcal{C}$. This means that $v_{H}$ can change the output labels of itself and its incident edges to get an improvement of at least $\lambda$. Let $\ell_{v}$ be that relabeling, so that we can get the improving set $(\set{v_{H}},\ell_{v})$, with $\operatorname{IR}(\set{v_{H}},\ell_{i},\ell_{v})\geq\lambda$. Let us now take the minimal version $(\set{v_{H}},\ell_{v}^{\star})$ of $(\set{v_{H}},\ell_{v})$.
Since $R_{c_{1}\operatorname{log}n}\leq\lambda$ by definition, Lemma 4.4 implies that $\set{v_{H}}$ is fully contained in
@@ 4.1 Technical claims and proofs
Here we prove Lemma 4.5 and Lemma 4.6. The first one basically says that, for any minimal improving set $A$ of some improving ratio $\beta$, for any node $v\in A$, we can find, within some radius-$O(\operatorname{log}n/\varepsilon)$ neighborhood of $v$, a minimal improving set $A^{\prime}\subseteq A$ with improving ratio at least $\beta-\varepsilon$. In order to prove it, we start with a preliminary lemma.
@@
We want to analyze the improvements of $(B,\ell_{B})$ and $(A,\ell_{A})$ compared to that of $(S,\ell^{\star}_{S})$. Consider any node $v$ such that its distance from $A$ (resp. $B$) is at least $r+1$. The two labelings $\ell_{A}$ (resp. $\ell_{B}$) and $\ell^{\star}_{S}$ will yield the same evaluation of the potential on the centered graph $(G_{r}(v),v)$.
@@
$(1+\varepsilon/(y(\Delta^{r}+1)))^{\left\lfloor(k-1)/2r\right\rfloor}$
@@
Now we want to prove Lemma 4.6. The key ingredient for Lemma 4.6 is the next lemma, which states the following: in a sequence of $\beta$-improving sets in $G$ w.r.t. some labeling $\ell$, for any node $v$ in any of the improving sets, there exists an improving set $(A,\ell_{A})$ in the radius-$O(\operatorname{log}^{2}n/\varepsilon^{2})$ neighborhood of $v$ (that is fully contained in the union of the improving sets) that has improving ratio at least $\beta-\varepsilon$ w.r.t. $\ell$. Now, we can take the minimal improving set $(A^{\prime},\ell_{A^{\prime}}^{\star})$ with $A^{\prime}\subseteq A$ and, by Lemma 4.5, we can find a set $A^{\prime\prime}\subseteq A^{\prime}$ such that $(A^{\prime\prime},\ell^{\star}_{A^{\prime\prime}})$ is a minimal improving set with improving ratio at least $\beta-2\varepsilon$ w.r.t. $\ell$ and diameter at most $O(\operatorname{log}n/\varepsilon)$. Hence, whenever there is a sequence of $\beta$-improving sets, before any relabeling, any node in the sequence can locally find small diameter improving sets with improving ratio at least $\beta-2\varepsilon$ w.r.t. $\ell$.
@@
Let $0<\varepsilon<\min\{\beta,\lambda\}$ be any value, for some $\beta>0$, where $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$. Let $((A_{i},\ell_{i}))_{1\leq i\leq h}$ be a sequence of $\beta$-improving sets w.r.t. $\ell$ in $G$, where each minimal improving set has diameter at most $c_{3}\operatorname{log}n/\varepsilon$ ($c_{3}$ as in Line 16). Then for all $1\leq j\leq h$, for all $v\in A_{j}$ there exists a large enough constant $c_{2}>0$ and some $k\leq c_{2}\operatorname{log}^{2}n/\varepsilon^{2}$, such that there exists a set $A\subseteq\mathcal{N}_{k}[v]\cap(\bigcup_{1\leq i\leq h}A_{i})$ with $\operatorname{IR}(A,\ell,\ell^{*}_{A})\geq\beta-\varepsilon$.
@@
Fix any $j\in[h]$, and consider an arbitrary $v\in A_{j}$. Let $x=\sigma(A_{j})\in V_{H}$. We define the “candidate sets” with regard to $z$ as $C_{i}=\mathcal{N}_{i}[z]$ for $i\geq 0$. Let $k$ be the minimum integer $i$ such that $\frac{\varepsilon}{K}w(C_{i})>w(C_{i+1}\setminus C_{i})=w(C_{i+1})-w(C_{i})$, for $K=(\beta-\varepsilon)+(\Delta^{r}+1)\Lambda$. We claim that $k\leq\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. By contradiction, suppose that $k>\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. Let $j=\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. Then, we have that $w(C_{i+1})-w(C_{i})\geq\frac{\varepsilon}{K}w(C_{i})$ for all $i\leq j$. Then,
@@
$w(C_{j})\geq\exp\left[\frac{KM\operatorname{log}(\Lambda n/\beta)}{\varepsilon+K}\right]\geq(\Lambda n/\beta)^{3/2}.$
@@
Note that $C_{k+1}=\mathcal{N}_{k+1}[z]$. Thus, the weak diameter of $\hat{A}$ in $G$ is at most $(2k+1)\cdot(2r+c_{3}\operatorname{log}n/\varepsilon)$. In particular, since $k\leq\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$ and $K\leq\beta+(\Delta^{r}+1)\Lambda$, we have that the weak diameter of $\hat{A}$ in $G$ is at most $100((\beta+\Delta^{r}\Lambda)\cdot\operatorname{log}(\Lambda n/\beta)/\varepsilon+1)(r+c_{3}\operatorname{log}n/\varepsilon)$. Since by assumption $\Delta$ and $r$ are constants, we have that $\hat{A}\subseteq\mathcal{N}_{c_{2}\operatorname{log}^{2}n/\varepsilon^{2}}[v]$ for some large enough constant $c_{2}>0$. By noticing that $\ell^{\star}_{\hat{A}}$ can only increase the improving ratio w.r.t. $\ell_{{A}}$, we conclude the proof of Lemma 4.8. ∎
@@
We first apply Lemma 4.8, and we obtain an improving set $(A,\ell_{A})$ with improving ratio at least $\beta-\varepsilon$ that is fully contained in $\mathcal{N}_{c_{2}\operatorname{log}^{2}/\varepsilon^{2}}[v]\cap(\cup_{i\in[h]}A_{i})$. Then, we take the minimal version of it, that is, $(A^{\prime},\ell^{\star}_{A^{\prime}})$, with $A^{\prime}\subseteq A$, and we apply Lemma 4.5, which yields the thesis. ∎

### v4

@@ 4 Analysis of the algorithm
This implies that there is a $j^{\star}\leq i-1$ such that $A$ is not contained in $\mathcal{N}_{t_{2}}[B_{j^{\star}}]$. Note that the diameter of $A$ is at most $c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon$ and at least one node of $A$ is not in $\mathcal{N}_{t_{2}}[B_{j^{\star}}]$. As a result, the set $\mathcal{N}_{c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda)/\varepsilon^{2}+2r+1}[A]$ is fully contained in some cluster $C^{(j^{\star})}_{s}$, for some $s\in[h_{j^{\star}}]$.
@@
$\mathcal{N}_{c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda)/\varepsilon^{2}}[v]\cap\left(\cup_{t\in[t_{i}]}\hat{S}_{t}\right),$
and such that $\operatorname{IR}(A^{\prime},\ell_{j^{\star}},\ell_{A^{\prime}})\geq R_{j^{\star}+1}-2\varepsilon$. By Lemma 4.5, for each $v\in A^{\prime}$, there exists a minimal improving set $(A^{\prime\prime},\ell_{A^{\prime\prime}})$ such that $A^{\prime\prime}\subseteq A^{\prime}\cap\mathcal{N}_{c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon}[v]$ and $\operatorname{IR}(A^{\prime\prime},\ell_{j^{\star}},\ell_{A^{\prime\prime}})\geq R_{j^{\star}+1}-3\varepsilon\geq R_{j^{\star}}$. The fact that
$A^{\prime\prime}\subseteq\mathcal{N}_{c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda)/\varepsilon^{2}}[v]$
implies that $A^{\prime\prime}\subseteq\mathcal{N}_{c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda)/\varepsilon^{2}}[A]$. Since $\mathcal{N}_{c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda)/\varepsilon^{2}+2r+1}[A]$ is fully contained in some cluster $C^{(j^{\star})}_{s}$, so it is $\mathcal{N}_{2r+1}[A^{\prime\prime}]$. Hence, the sequence $(S^{(j^{\star})}_{s,t},\ell_{S^{(j^{\star})}_{s,t}})_{t\in[k_{s}^{(j^{\star})}]}$ was not maximal, reaching a contradiction. ∎
@@
By Lemma 4.2, we have that the algorithm terminates. As discussed in Section 3.2, the running time of the algorithm is $O(\Delta^{2r}(\Lambda/\lambda)^{2}\operatorname{log}^{6}(\Lambda n/\lambda))$. We just have to argue about correctness.
Let $i=c_{1}\operatorname{log}n$ be the last phase, and let $\ell_{i}$ be the labeling obtained at the end of phase $i$ (Lines 18 and 23). By contradiction, suppose that there is an error in the graph, that is, a centered graph $(H,v_{H})$ of radius $2r$ that does not belong to $\mathcal{C}$. This means that $v_{H}$ can change the output labels of itself and its incident edges to get an improvement of at least $\lambda$. Let $\ell_{v}$ be that relabeling, so that we can get the improving set $(\{v_{H}\},\ell_{v})$, with $\operatorname{IR}(\{v_{H}\},\ell_{i},\ell_{v})\geq\lambda$. Let us now take the minimal version $(\{v_{H}\},\ell_{v}^{\star})$ of $(\{v_{H}\},\ell_{v})$.
Since $R_{c_{1}\operatorname{log}n}\leq\lambda$ by definition, Lemma 4.4 implies that $\{v_{H}\}$ is fully contained in
@@ 4.1 Technical claims and proofs
Here we prove Lemma 4.5 and Lemma 4.6. The first one basically says that, for any minimal improving set $A$ of some improving ratio $\beta$, for any node $v\in A$, we can find, within some radius-$O(\Delta^{r}\Lambda\operatorname{log}n/\varepsilon)$ neighborhood of $v$, a minimal improving set $A^{\prime}\subseteq A$ with improving ratio at least $\beta-\varepsilon$. In order to prove it, we start with a preliminary lemma.
@@
We want to analyze the improvements of $(B,\ell_{B})$ and $(A,\ell_{A})$ compared to that of $(S,\ell^{\star}_{S})$. Consider any node $v$ such that its distance from $A$ (resp. $B$) is at least $r+1$. The two labelings $\ell_{B}$ (resp. $\ell_{A}$) and $\ell^{\star}_{S}$ will yield the same evaluation of the potential on the centered graph $(G_{r}(v),v)$.
@@
$\left(1+\frac{\varepsilon}{y(\Delta^{r}+1)}\right)^{\left\lfloor(k-1)/2r\right\rfloor}$
@@
Now we want to prove Lemma 4.6. The key ingredient for Lemma 4.6 is the next lemma, which states the following: in a sequence of $\beta$-improving sets in $G$ w.r.t. some labeling $\ell$, for any node $v$ in any of the improving sets, there exists an improving set $(A,\ell_{A})$ in the radius-$O(\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda)/\varepsilon^{2})$ neighborhood of $v$ (that is fully contained in the union of the improving sets) that has improving ratio at least $\beta-\varepsilon$ w.r.t. $\ell$. Now, we can take the minimal improving set $(A^{\prime},\ell_{A^{\prime}}^{\star})$ with $A^{\prime}\subseteq A$ and, by Lemma 4.5, we can find a set $A^{\prime\prime}\subseteq A^{\prime}$ such that $(A^{\prime\prime},\ell^{\star}_{A^{\prime\prime}})$ is a minimal improving set with improving ratio at least $\beta-2\varepsilon$ w.r.t. $\ell$ and diameter at most $O(\Delta^{r}\Lambda\operatorname{log}n/\varepsilon)$. Hence, whenever there is a sequence of $\beta$-improving sets, before any relabeling, any node in the sequence can locally find small diameter improving sets with improving ratio at least $\beta-2\varepsilon$ w.r.t. $\ell$.
@@
Let $0<\varepsilon<\lambda$ be any value, where $\lambda$ is the minimum improvement associated to $(\Pi,\Psi)$, and let $\beta\geq\lambda$. Let $((A_{i},\ell_{i}))_{1\leq i\leq h}$ be a sequence of $\beta$-improving sets w.r.t. $\ell$ in $G$, where each minimal improving set has diameter at most $c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon$ ($c_{3}$ as in Line 16). Then for all $1\leq j\leq h$, for all $v\in A_{j}$ there exists a large enough constant $c_{2}>0$ and some $k\leq c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda)/\varepsilon^{2}$, such that there exists a set $A\subseteq\mathcal{N}_{k}[v]\cap(\bigcup_{1\leq i\leq h}A_{i})$ with $\operatorname{IR}(A,\ell,\ell^{*}_{A})\geq\beta-\varepsilon$.
@@
Fix any $j\in[h]$, and consider an arbitrary $v\in A_{j}$. Let $z=\sigma(A_{j})\in V_{H}$. We define the “candidate sets” with regard to $z$ as $C_{i}=\mathcal{N}_{i}[z]$ for $i\geq 0$. Let $k$ be the minimum integer $i$ such that $\frac{\varepsilon}{K}w(C_{i})>w(C_{i+1}\setminus C_{i})=w(C_{i+1})-w(C_{i})$, for $K=(\beta-\varepsilon)+(\Delta^{r}+1)\Lambda$. We claim that $k\leq\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. By contradiction, suppose that $k>\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. Let $j=\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$. Then, we have that $w(C_{i+1})-w(C_{i})\geq\frac{\varepsilon}{K}w(C_{i})$ for all $i\leq j$. Then,
@@
$w(C_{j})\geq\exp\left[\frac{3K\operatorname{log}(\Lambda n/\beta)}{\varepsilon+K}\right]\geq(\Lambda n/\beta)^{3/2}.$
@@
Note that $C_{k+1}=\mathcal{N}_{k+1}[z]$. Thus, the weak diameter of $\hat{A}$ in $G$ is at most $(2k+1)\cdot(2r+c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon)$. In particular, since $k\leq\left\lceil 3K\operatorname{log}(\Lambda n/\beta)/\varepsilon\right\rceil$ and $K\leq\beta+(\Delta^{r}+1)\Lambda$, we have that the weak diameter of $\hat{A}$ in $G$ is at most $100((\beta+\Delta^{r}\Lambda)\cdot\operatorname{log}(\Lambda n/\beta)/\varepsilon+1)(r+c_{3}\Delta^{r}\Lambda\operatorname{log}n/\varepsilon)$. Since by assumption $r$ is constant and $\lambda\leq\beta\leq\Lambda$, we have that
$\hat{A}\subseteq\mathcal{N}_{c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda)/\varepsilon^{2}}[v],$
for some large enough constant $c_{2}>0$. By noticing that $\ell^{\star}_{\hat{A}}$ can only increase the improving ratio w.r.t. $\ell_{{A}}$, we conclude the proof of Lemma 4.8. ∎
@@
We first apply Lemma 4.8, and we obtain an improving set $(A,\ell_{A})$ with improving ratio at least $\beta-\varepsilon$ that is fully contained in $\mathcal{N}_{c_{2}\Delta^{2r}\Lambda^{2}\operatorname{log}^{2}(\Lambda n/\lambda)/\varepsilon^{2}}[v]\cap(\cup_{i\in[h]}A_{i})$. Then, we take the minimal version of it, that is, $(A^{\prime},\ell^{\star}_{A^{\prime}})$, with $A^{\prime}\subseteq A$, and we apply Lemma 4.5, which yields the thesis. ∎
5 A lower bound for locally optimal cut
In this section we prove a lower bound of $\Omega(\min\{\Delta,\sqrt{n}\})$ rounds for any algorithm that finds a locally optimal cut. Naturally this implies that any algorithm for generic local potential problems must also take at least $\Omega(\min\{\Delta,\sqrt{n}\})$ rounds.
Definition 5.1 (Locally optimal cut).
Let $G=(V,E)$ be a graph. A locally optimal cut is an assignment of labels $\ell:V\to\{-1,1\}$ such that, for every node $v\in V$, the number of neighbors of $v$ with a label different from $\ell(v)$ is at least as large as the number of neighbors that share $\ell(v)$, i.e.,
[…]
We prove the following lower bound.
Theorem 5.2.
For any large enough $n\in\mathbb{N}$ and $\Delta<n$, let $\varepsilon=1-(\Delta/n)$. Any algorithm that finds a locally optimal cut in the randomized LOCAL model with probability $p\geq\frac{1}{2}$ on graphs of $n$ nodes and maximum degree $\Delta$ has locality $\Omega(\min\{\Delta,\sqrt{\varepsilon n}\})$. In particular, for any $\Delta\leq cn$ for some constant $c<1$, the locality is $\Omega(\min\{\Delta,\sqrt{n}\})$.
The proof idea is the following: We construct a graph that contains a component of diameter $\Theta(\min\{\Delta,\sqrt{\varepsilon n}\})$. Such component contains two nodes $u$ and $v$ at distance $\Theta(\min\{\Delta,\sqrt{\varepsilon n}\})$ satisfying the following property: in any valid solution to the locally optimal cut problem, the nodes $u$ and $v$ must have different labels.
The following lemma will be a key ingredient in our lower bound construction.
Lemma 5.3.
Let $G=(V,E)$ be a graph, and let $u,v\in V$ be two nodes of odd degree that share the same open neighborhood, i.e., $\mathcal{N}(u)=\mathcal{N}(v)$. Then, in any valid solution to the locally optimal cut problem, it must hold that $u$ and $v$ have the same label.
Proof.
Let $\ell$ be any valid labeling. For a node $x\in V$, denote by
$P_{x}=\{\,y\in\mathcal{N}(x)\mid\ell(y)=1\,\}\quad\text{and}\quad N_{x}=\{\,y\in\mathcal{N}(x)\mid\ell(y)=-1\,\}$
the sets of neighbors of $x$ labeled $1$ and 1-1, respectively. Since $u$ has odd degree, exactly one of the sets $P_{u}$ or $N_{u}$ must be larger. By the definition of a locally optimal cut, node $u$ must be labeled $1$ if $|N_{u}|>|P_{u}|$, and 1-1 if $|P_{u}|>|N_{u}|$. The same reasoning applies to $v$. Because $\mathcal{N}(u)=\mathcal{N}(v)$, we have $P_{u}=P_{v}$ and $N_{u}=N_{v}$. Hence, both $u$ and $v$ must have the same label. ∎
5.1 The lower bound graph
We build our lower bound graphs by merging two separate copies of the following construction. For an integer parameter $k\geq 2$, we define a graph $H_{k}=(V,E)$ as follows.
Node set.
The vertex set $V$ is partitioned into disjoint subsets, which we call layers, arranged in the sequence
$S,L_{2},R_{2},L_{3},R_{3},\ldots,L_{k},R_{k}.$
Here, $S$ is a singleton set ($|S|=1$), and for each $i\in\{2,\ldots,k\}$, the sets $L_{i}$ and $R_{i}$ have cardinality $|L_{i}|=|R_{i}|=i$.
Edge set.
The edge set $E$ is defined so that each pair of consecutive sets in the sequence
$S,L_{2},R_{2},L_{3},R_{3},\ldots,L_{k},R_{k}$
forms a complete bipartite graph. Namely, for each $i\in\{2,\ldots,k\}$, all nodes in $L_{i}$ are connected to all nodes in $R_{i}$, and all nodes in $R_{i}$ are connected to all nodes in $L_{i+1}$ (if $i<k$). As for the only node in $S$, it is connected to all nodes in $L_{2}$.
Hence, the sequence of layers induces the following pattern of complete bipartite graphs:
$K_{1,2},\;K_{2,2},\;K_{2,3},\;K_{3,3},\;K_{3,4},\;\ldots,\;K_{k-1,k-1},\;K_{k-1,k},\;K_{k,k}.$
Properties.
Intuitively, the graph can be seen as a “path” of layers whose sizes increase monotonically as
$1,2,2,3,3,4,4,\ldots,k-1,k-1,k,k,$
where adjacent layers are fully connected. Note that all nodes in a specific layer, except the layer $S$ and (possibly) the last layer, have an odd degree and share the same neighborhood. The following lemma shows that the labels of nodes in any given layer are all the same, while the labels of nodes of consecutive layers must differ.
Lemma 5.4.
Consider $H_{k}$ to be the input graph. In any solution to the locally optimal cut problem, all nodes in any layer of $H_{k}$ share the same label, while nodes in consecutive layers of $H_{k}$ use different labels.
Proof.
Any node in any layer $L_{i}$, $2\leq i\leq k$, and in any layer $R_{j}$, $2\leq j\leq k-1$, has odd degree. Hence, by Lemma 5.3, all nodes in $L_{i}$ must share the same label for $2\leq i\leq k$ (the same holds in $R_{j}$ for $2\leq j\leq k-1$), implying that also nodes in $R_{k}$ must share the same label.
We now prove the second part of the statement. It is straightforward to notice that the node in $S$ needs to pick a label different from the nodes in $L_{2}$. Similarly, the nodes in $R_{k}$ must have the opposite label of the nodes in $L_{k}$. Now fix any three consecutive layers in the sequence
$S,L_{2},R_{2},L_{3},R_{3},\ldots,L_{k},R_{k}$
We denote these consecutive layers as $A_{i-1}$, $A_{i}$, and $A_{i+1}$. Each node in $A_{i}$ is adjacent to all nodes in $A_{i-1}$ and $A_{i+1}$, and to no other nodes. Because $|A_{i+1}|>|A_{i-1}|$, the majority of $A_{i}$’s neighbors are in $A_{i+1}$.
By contradiction, assume that the nodes in $A_{i}$ and the nodes in $A_{i+1}$ use the same label and, w.l.o.g., let that label be +1+1. Then every node in $A_{i}$ has strictly more neighbors with label +1+1 (from $A_{i+1}$) than it could have with label 1-1 (from $A_{i-1}$). This violates the locally optimal cut condition and, hence, the nodes in $A_{i}$ and those in $A_{i+1}$ cannot use the same label. We just proved that nodes in consecutive layers must be labeled differently. ∎
We now complete the lower bound construction by “mirroring” the graph $H_{k}$.
Mirrored construction.
We consider a graph $G_{k}$ constructed as follows. Consider a copy $H_{k}^{(1)}$ of $H_{k}$ and a copy $H_{k-1}^{(2)}$ of $H_{k-1}$, where
$S^{(1)},L_{2}^{(1)},R_{2}^{(1)},L_{3}^{(1)},R_{3}^{(1)},\ldots,L_{k}^{(1)},R_{k}^{(1)}$
are the layers of $H_{k}^{(1)}$, and
$S^{(2)},L_{2}^{(2)},R_{2}^{(2)},L_{3}^{(2)},R_{3}^{(2)},\ldots,L_{k-1}^{(2)},R_{k-1}^{(2)}$
are the layers of $H_{k-1}^{(2)}$. Now, we connect nodes in $R_{k}^{(1)}$ to all nodes in $R_{k-1}^{(2)}$ to form a complete bipartite graph. The structure of layers of $G_{k}$ is then
$S^{(1)},L_{2}^{(1)},R_{2}^{(1)},L_{3}^{(1)},R_{3}^{(1)},\ldots,L_{k}^{(1)},R_{k}^{(1)},R_{k-1}^{(2)},L_{k-1}^{(2)},\ldots,R_{3}^{(2)},L_{3}^{(2)},R_{2}^{(2)},L_{2}^{(2)},S^{(2)},$
and the sizes of layers are as follows:
$1,2,2,3,3,4,4,\ldots,k-1,k-1,k,k,k-1,k-1,\ldots,3,3,2,2,1.$
Note that the nodes in $S^{(1)},S^{(2)}$ are at odd distance. If we prove that, with this new construction, the layers $L_{k}^{(1)},R_{k}^{(1)},R_{k-1}^{(2)},L_{k-1}^{(2)}$ are still monochromatic and must alternate colors, then we are asking the nodes in $S^{(1)},S^{(2)}$ to have different labels.
Lemma 5.5.
For any $k\geq 2$, the graph $G_{k}$ has the following properties:
  1. (i)
    The number of nodes in $G_{k}$ is $2k^{2}-2$.
  2. (ii)
    The maximum degree of $G_{k}$ is 2k12k-1.
  3. (iii)
    In any correct solution to the locally optimal cut problem, each layer is monochromatic and adjacent layers use different labels.
Proof.
Consider claim (i). There are 2 layers with 1 node (the first and last), 2 layers with $k$ nodes (in the middle) and 4 layers of any other cardinality in $\{2,\ldots,k-1\}$. Therefore, the following holds:
[…]
We are ready to prove Theorem 5.2 through an indistinguishability argument.
Proof of Theorem 5.2.
Let $\mathcal{A}$ be a $T(\Delta,n)$-round randomized LOCAL algorithm that solves locally optimal cut on graphs of $n$ nodes and maximum degree $\Delta$, with $\Delta\leq n(1-\varepsilon)$. We will show that, for all large enough values of $n$, and for all large enough values of $\Delta\leq n(1-\varepsilon)$, if $\mathcal{A}$ has success probability at least $1/2$, then it must hold that $T(\Delta,n)>\left\lfloor\min\{(\Delta+1)/2,\sqrt{\varepsilon n}/2\}\right\rfloor/2$.
[…]
It is easy to see that all the properties of Lemma 5.5 still hold for the induced subgraph $G[V(G_{k})]$, and, in particular, by Property (iii), $u$ and $v$ must output different labels in any valid solution to the locally optimal cut problem on $G$.
By contradiction, assume that $T(\Delta,n)\leq\left\lfloor\min\{(\Delta+1)/2,\sqrt{\varepsilon n}/2\}\right\rfloor/2$ (which is less than $k/2$), and that $\mathcal{A}$ succeeds with probability at least $1/2$. By construction of $G$, the radius-$T$ views of $u$ and $v$ are isomorphic, since $T=T(n,\Delta)\leq k/2=\Theta(\sqrt{n})$ and both $P_{1}$ and $P_{2}$ have length $\Theta(n)$. Hence, by indistinguishability, there is a probability of at least $1/2$ that both nodes in $S^{(1)}$ and $S^{(2)}$ take the same output label, reaching a contradiction. ∎
In Appendix B, we show that the same method can be applied to prove a lower bound of $\Omega(\min\{\Delta,\sqrt{n}\})$ rounds for the locally optimal cut problem even in the quantum-LOCAL model. The following is a corollary of Theorem B.1.
Corollary 5.6.
For any large enough $n\in\mathbb{N}$ and $\Delta<n$, let $\varepsilon=1-(\Delta/n)$. Any algorithm that finds a locally optimal cut in the quantum-LOCAL model with probability $p\geq\frac{1}{2}$ on graphs of $n$ nodes and maximum degree $\Delta$ has locality $\Omega(\min\{\Delta,\sqrt{\varepsilon n}\})$. In particular, for any $\Delta\leq cn$ for some constant $c<1$, the locality is $\Omega(\min\{\Delta,\sqrt{n}\})$.